Find the electric flux through a cylinder with a distance of 24 cm and a height of 37 cm with a charge of 5.42 μC. ( use the general formula of flux and the area of cylinder is 2𝜋rh + 2𝜋r²)



1
Expert's answer
2022-03-18T09:08:06-0400

The electric flux is given by

"\Phi=\oint EdA=\frac{q}{\epsilon_0}"

"\Phi=\frac{5.42\times 10^{-6}}{8.85\times 10^{-12}}=6.12\times 10^5\:\rm N\cdot m^2/C"